# DeformTime: Capturing Variable Dependencies with Deformable Attention for Time Series Forecasting

### v2

@@ 5 Conclusions
We propose DeformTime, a novel deep learning architecture for multivariate time series forecasting that uses deformable attention blocks to effectively learn from exogenous predictors, deploying specific operations that aim to capture inter- and intra-variable dependencies. We assess forecasting accuracy using $3$ established benchmark tasks as well as $3$ influenza-like illness prediction tasks with web search frequency time series as exogenous variables, covering $3$ locations (in $2$ countries) and spanning a time period longer than $12$ years. DeformTime yields strong forecasting accuracy across the board, reducing mean absolute error on average by $10\%$ compared to the best baseline model (different for each task and horizon) and by at least $15.2\%$ (for Crossformer [55]) when compared to a specific model. Importantly, performance gains remain stable for longer forecasting horizons. Our experiments, including the ablation study, highlight that modelling variable dependencies is an important attribute. Specifically in the disease forecasting task, where more exogenous predictors are present, the most competitive baselines capture variable dependencies to some extent, whereas models that do not on many occasions cannot surpass the performance of a naïve persistence model (see also Appendix D.7). In contrast to other methods, DeformTime’s GPU memory footprint is not significantly affected by the amount of exogenous variables. We discuss the limitations, societal impact, and future directions for this work in Appendices E and F.
@@ Acknowledgements
V.L. acknowledges all levels of support from the EPSRC grant EP/X031276/1 and a Google donation.
Supplementary Materials
\@bottomtitlebar
@@ Appendix A Data sets for time series forecasting
This section offers a detailed description of the data sets used in our experiments to assess the performance of a series of methods for MTS forecasting. We provide links to data sets that are publicly available. Data sets that were formed by us will be provided in our source code repository. If we cannot share the exact data set due to copyright restrictions, we will provide instructions as to how research teams can obtain access directly from the data provider.
A.1 Established MTS forecasting benchmarks
@@
ETTh1 and ETTh2 are electricity transformer temperature data sets that were obtained from $2$ different counties in China [57]. [^8] ETTh1 and ETTh2, github.com/zhouhaoyi/ETDataset Each data set contains $6$ exogenous variables capturing power load attributes, and the target variable which is the temperature of oil. The data set covers a period from July 1, 2016 to June 26, 2018 (although not all data points are used in our experiments to match the setup in related methods that we compare to). The temporal resolution of these data sets is hourly. Adopting the setup in prior work [46, 32, 28], we use a total amount of $14{,}400$ time steps (starting from July 1, 2016) where the first $8{,}640$ are used for training, the next $2{,}880$ for validation, and the last $2{,}880$ for testing.
@@
The weather data set [45] contains meteorological measurements collected from a weather station at the Max Planck Biogeochemistry Institute. It covers a year from January 1, 2020 to January 1, 2021.999The data set can be obtained from Autoformer’s repository, github.com/thuml/Autoformer The temporal resolution is $10$ minutes and $20$ meteorological indicators (exogenous variables) are being reported; the target variable is wet-bulb temperature. This results in a total of $52{,}696$ time steps (samples). Based on prior work [55], we use $70\%$, $10\%$, and $20\%$ of the time steps for training, validation, and testing, respectively.
@@ A.2 Forecasting influenza-like illness rates using web search activity
Important components of our data set are shared in the paper’s online repository.1 The shared information allows replication of our work for teams who have obtained the same data-sharing agreements (mainly access to the Google Health Trends API). However, we anticipate that we will be permitted to share all data sets (to make replication and comparisons easier).
@@
We also note that in practice, the reported ILI rate is delayed, i.e. assuming that it is published by RCGP or CDC at time $t$, it actually refers to an ILI rate representing a previous time $t\!-\!\delta$. Specifically, we consider that there is a delay of $\delta\!=\!7$ days for RCGP/England, and a delay of $\delta\!=\!14$ days for CDC/US Regions [42, 36]. This delay, $\delta$, has an impact on the autoregressive time series (of the target variable, in this case, the ILI rate) in a forecasting task (see also section 2). As the delay increases, the forecaster is expected to rely more on the exogenous predictors (that are not delayed) rather than the autoregressive time series of the target variable.
@@
Web search activity trends, if modelled appropriately [10, 22, 20, 18], are a good indicator of influenza rates in a population and can become a strong exogenous predictor for influenza forecasting [9, 31]. We obtain web search activity data from the Google Health Trends API; this is not a publicly available API, but access can be provided via an application process. [^11] Google Health Trends API access application, support.google.com/trends/contact/trends_api For a day and a certain location, the frequency of a search query is determined as the ratio of searches conducted for a particular term or set of terms divided by the overall search volume. We use a pool of $22{,}071$ unique health-related search queries and obtain their daily frequency from August 1 (US regions) or September 1 (England), 2006 to July 31 (US regions) or August 31 (England), 2019. However, we note that not all search queries are used as exogenous variables in our forecasting models. A feature selection process is described in Appendix A.3.
@@
We assess the accuracy of forecasters across the last $4$ influenza seasons ($4$ separate test sets) in England and the US regions. Each test period is a complete influenza season (as previously defined). For each test period, we train models based on the $9$ influenza seasons that precede it. A part of each training set is used for validating model decisions, including hyperparameters. We construct validation sets using the following strategy. Each validation set has $180$ days in total, using $60$-day periods to capture the onset, peak, and the period after the peak (or the outset) of an influenza season. Each $60$-day period comes from a different influenza season in the training set; we use the last $3$ influenza seasons in the training data to make sure our validation process has a recency effect. The last, penultimate, and third from last influenza seasons are used to define the outset, peak, and onset validation periods, respectively. We use CDC’s definition to determine the onset of an influenza season, i.e. a time point is deemed to be the onset when the subsequent ILI rates exceed a threshold for $2$ consecutive weeks. [^12] CDC’s influenza season onset definition, cdc.gov/flu/weekly/overview.htm The peak point is simply the highest ILI rate within an influenza season. The outset has the inverse definition from the onset, i.e. the last time point where the ILI rate exceeds the onset threshold for two consecutive weeks. Once these time points are determined, they become the $30$th point in a $60$-day validation period. Figure A1 shows an example of the validation periods that were determined when training an ILI rate model for England using (training) data from September 2009 to August 2018 (with the test period being September 2018 to August 2019).
Figure A1: An example of how the training, validation, and test sets are constructed when the test influenza season is 2018/19 (England). The lines in blue, red, and orange colour denote the training, validation, and test periods, respectively. To form the validation set from our training data, we select the period after the peak (outset) from the third to last influenza season, the period around the peak from the penultimate season, and a period around influenza onset from the last season.
@@ A.3 Feature selection for web search activity time series
Originally, we consider $22{,}071$ search queries. We then perform two feature selection steps to maintain more relevant queries to the ILI forecasting task. First, we apply a semantic filter to remove queries that are not related to the topic of influenza, similarly to the approach presented by Lampos et al. [21]. We obtain an embedding representation for each search query using a pre-trained sentence BERT model. [^13] Sentence BERT, huggingface.co/bert-base-uncased We also obtain the embeddings of influenza-related (base) terms and expressions, such as “flu”, “flu symptoms” etc. We compute the cosine similarity between each base term and search query and maintain the top-$1000$ (England) or top-$500$ (US regions) search queries (per base term). This leaves $4{,}396$ and $2{,}479$ queries respectively for England and the US regions. [^14] Base and filtered search terms are provided in DeformTime’s repository (see footnote 1). The second selection step is dynamic, i.e. it might have a different outcome as the training data change. In this step, we compute the linear correlation between the remaining search queries and the target variable (ILI rate) in the $5$ most recent influenza seasons in the training data (to impose recency). We maintain search queries with a correlation that exceeds a correlation threshold $\tau$, which is a learnable hyperparameter (see also Appendix C).
A.4 Data normalisation
In all forecasting tasks, we standardise all variables, an operation also known as z-scoring (zero mean, unit standard deviation), each time based on the training data. The prediction output is de-normalised back to its original scale prior to being compared with the (de-normalised) ground truth. In addition to that, for the ETT and weather data sets, we also z-score each variable within the input’s look-back window, in accordance with previous work [27, 58, 32].
A.5 Criticism of data sets used to benchmark forecasting methods
@@
Although we present results using the ETTh1, ETTh2, and weather data sets to compare with other forecasting methods in the literature, we consider these data sets to have several shortcomings. Clearly, a weather data set that covers only a $1$-year period and is based on the sensors from just one geographical location cannot be expected to provide good enough insights into meteorological forecasting. It captures a negligible portion of meteorological data and, as a result, derived forecasting models cannot have any practical impact (compared to actual models that are used in meteorology). Hence, using this data set for training and evaluating a forecasting model may, unfortunately, result in misleading or at least inconclusive outcomes. Similar issues, but perhaps to a smaller extent, are present in the ETTh1 and ETTh2 data sets. On the one hand, these data sets seem to have a more solid practical application. However, experiments on their ETTm1 and ETTm2 variants, where the temporal resolution changes from $1$ hour to $15$ minutes (same data, shorter temporal resolution), showcase part of the problem. When we compared the forecasting performance of SOTA baselines between the two different temporal resolutions, we surprisingly found that the prediction based on the hourly sampled data set has better accuracy compared to the quarter-hourly sampled data set even at longer forecasting horizons. For example, models that conduct $192$ or even $336$ time steps (equivalent to $8$ or $14$ days) ahead forecasting with ETTh1 and ETTh2 achieve a lower MAE compared to a $720$ time steps ahead forecast with ETTm1 and ETTm2 (equivalent to $7.5$ days ahead forecast), respectively [32, 46, 29].
@@
We also noticed that some papers run experiments on US national weekly ILI rates obtained from the CDC [59, 43, 32]. The first concerning observation was that in these experiments, the input variables were based on variates (columns) from the CDC extract (spreadsheet) that only provide redundant insights compared to the ILI rate (e.g. the raw numerator and denominator of an ILI rate or the population unweighted ILI rate). In addition, some forecasting horizons that were explored in these benchmarks were practically unreasonable (even from a biological perspective, there are limits in predicting the future, especially when it comes to viruses). SOTA forecasting models for influenza generally provide good accuracy at approximately a $2$ weeks ahead forecasting horizon [33, 31]. Anything beyond that with satisfactory performance should be considered as a very important development within the disease modelling community. We do think DeformTime might fall in that category as it delivers good results for $3$ or $4$ weeks ahead forecasting horizons. Contrary to that, in the aforementioned ILI rate benchmark, there were forecasting horizons of $24$ or $60$ weeks ahead (i.e. even more than a year!). Of course, the accuracy of the forecasting models under these forecasting horizons was reportedly detrimentally poor (influenza does not have a strong periodicity), and consequently could not be used for any meaningful comparative conclusions. If all forecasting models are performing poorly, even the least poorly performing one still is an inadequate forecaster (for the underlying task).
@@
Beyond this, we have also excluded data sets, such as traffic [45] and electricity consumption [57], that do not have one explicitly defined target variable, but instead focus on multivariate predictions (multi-task learning). Consequently, these data sets are not entirely compatible with our forecasting task or method (DeformTime makes predictions about one target variable).
@@
Appendix B Summary of other forecasting models

### v3

@@ 5 Conclusions
We propose DeformTime, a novel deep learning architecture for multi-variable time series forecasting that uses deformable attention blocks to effectively learn from exogenous predictors, deploying specific operations that aim to capture inter- and intra-variable dependencies. We assess forecasting accuracy using $3$ established benchmark tasks as well as $3$ ILI rate prediction tasks with web search frequency time series as exogenous variables, covering $3$ locations (in $2$ countries) and spanning a time period longer than $12$ years. DeformTime yields strong forecasting accuracy across the board, reducing mean absolute error on average by $7.2\%$ compared to the best baseline model (different for each task and horizon) and by at least $11.9\%$ (for ModernTCN (Luo & Wang, 2024b)) when compared to a specific model. Importantly, performance gains remain stable for longer forecasting horizons. Our experiments, including the ablation study, highlight that modelling variable dependencies is an important attribute. Specifically in the disease forecasting task, where more exogenous predictors are present, the most competitive baselines capture variable dependencies to some extent, whereas models that do not on many occasions cannot surpass the performance of a naïve persistence model (see also Appendix E.9). In contrast to other methods, DeformTime’s GPU memory footprint is not significantly affected by the amount of exogenous variables.
@@ Acknowledgements
The authors would like to thank the RCGP for providing ILI rates for England. V. Lampos would like to acknowledge all levels of support from the EPSRC grant EP/X031276/1.
Supplementary Material / Appendix
@@ Appendix A Data sets for time series forecasting
This section offers a detailed description of the data sets used in our experiments to assess the performance of a series of methods for MTS forecasting. We provide links to data sets that are publicly available. For the ILI data sets formed by us, while we hope that we will be able to share the exact data set with the broader community in the future, we cannot do so due to copyright restrictions at the time of publication. As an alternative, we provide instructions as to how research teams can obtain access directly from the data provider.
A.1 Established time series forecasting benchmarks
@@
ETTh1 and ETTh2 are electricity transformer temperature data sets that were obtained from $2$ different counties in China (Zhou et al., 2021). [^10] ETTh1 and ETTh2, github.com/zhouhaoyi/ETDataset Each data set contains $6$ exogenous variables capturing power load attributes, and the target variable which is the temperature of oil. The data set covers a period from July 1, 2016 to June 26, 2018 (although not all data points are used in our experiments to match the setup in related methods that we compare to). The temporal resolution of these data sets is hourly. Adopting the setup in prior work (Wu et al., 2023; Nie et al., 2023; Liu et al., 2024), we use a total amount of $14{,}400$ time steps (starting from July 1, 2016) where the first $8{,}640$ are used for training, the next $2{,}880$ for validation, and the last $2{,}880$ for testing.
@@
The weather data set (Wu et al., 2021) contains meteorological measurements collected from a weather station at the Max Planck Biogeochemistry Institute. It covers a year from January 1, 2020 to January 1, 2021.111111Weather data set, github.com/thuml/Autoformer The temporal resolution is $10$ minutes and $20$ meteorological indicators (exogenous variables) are being reported; the target variable is carbon dioxide concentration level. This results in a total of $52{,}696$ time steps (samples). Based on prior work (Zhang & Yan, 2023), we use $70\%$, $10\%$, and $20\%$ of the time steps for training, validation, and testing, respectively.
@@ A.2 Forecasting influenza-like illness rates using web search activity
Important components of our data set are shared in the paper’s online repository. At the time of publication, these allow replication of our work for teams who have obtained the same data-sharing agreements (mainly access to the Google Health Trends API).
@@
We also note that in practice, the reported ILI rate is delayed, i.e. assuming that it is published by RCGP or CDC at time $t$, it actually refers to an ILI rate representing a previous time $t\!-\!\delta$. Specifically, we consider that there is a delay of $\delta\!=\!7$ days for RCGP/England, and a delay of $\delta\!=\!14$ days for CDC/US Regions (Wagner et al., 2018; Reich et al., 2019). This delay, $\delta$, has an impact on the autoregressive time series (of the target variable, in this case, the ILI rate) in a forecasting task (see also section 2). As the delay increases, the forecaster is expected to rely more on the exogenous predictors (that are not delayed) rather than the autoregressive time series of the target variable.
@@
Web search activity trends, if modelled appropriately (Ginsberg et al., 2009; Lazer et al., 2014; Lampos et al., 2015; Kandula & Shaman, 2019), are a good indicator of influenza rates in a population and can become a strong exogenous predictor for influenza forecasting (Dugas et al., 2013; Morris et al., 2023). We obtain web search activity data from the Google Health Trends API; this is not a publicly available API, but access can be provided via an application process. [^13] Google Health Trends API, support.google.com/trends/contact/trends_api For a day and a certain location, the frequency of a search query is determined as the ratio of searches conducted for a particular term or set of terms divided by the overall search volume. We use a pool of $22{,}071$ unique health-related search queries and obtain their daily frequency from August 1 (US regions) or September 1 (England), 2006 to July 31 (US regions) or August 31 (England), 2019. However, we note that not all search queries are used as exogenous variables in our forecasting models. A feature selection process is described in Appendix A.3.
@@
Figure S1: An example of how the training, validation, and test sets are constructed when the test influenza season is 2018/19 (England). The lines in blue, red, and orange colour denote the training, validation, and test periods, respectively. To form the validation set from our training data, we select the period after the peak (outset) from the third to last influenza season, the period around the peak from the penultimate season, and a period around influenza onset from the last season.
@@
We assess the accuracy of forecasting models across the last $4$ influenza seasons ($4$ separate test sets) in England and the US regions. Each test period is a complete influenza season (as previously defined). For each test period, we train models based on the $9$ influenza seasons that precede it. A part of each training set is used for validating model decisions, including hyperparameters. We construct validation sets using the following strategy. Each validation set has $180$ days in total, using $60$-day periods to capture the onset, peak, and the period after the peak (or the outset) of an influenza season. Each $60$-day period comes from a different influenza season in the training set; we use the last $3$ influenza seasons in the training data to make sure our validation process has a recency effect. The last, penultimate, and third from last influenza seasons are used to define the outset, peak, and onset validation periods, respectively. We use CDC’s definition to determine the onset of an influenza season, i.e. a time point is deemed to be the onset when the subsequent ILI rates exceed a threshold for $2$ consecutive weeks. [^14] CDC’s influenza season onset definition, cdc.gov/fluview/overview The peak point is simply the highest ILI rate within an influenza season. The outset has the inverse definition from the onset, i.e. the last time point where the ILI rate exceeds the onset threshold for two consecutive weeks. Once these time points are determined, they become the $30$th point in a $60$-day validation period. Figure S1 shows an example of the validation periods that were determined when training an ILI rate model for England using (training) data from September 2009 to August 2018 (the test period being September 2018 to August 2019).
@@ A.3 Feature selection for web search activity time series
Originally, we consider $22{,}071$ search queries. We then perform two feature selection steps to maintain more relevant queries to the ILI forecasting task. First, we apply a semantic filter to remove queries that are not related to the topic of influenza, similarly to the approach presented by Lampos et al. (2017). We obtain an embedding representation for each search query using a pre-trained sentence BERT model. [^15] Sentence BERT, huggingface.co/bert-base-uncased We also obtain the embeddings of influenza-related (base) terms and expressions, such as “flu”, “flu symptoms” and so on. [^16] The complete list of base terms is provided at github.com/ClaudiaShu/DeformTime. We compute the cosine similarity between each base term and search query and maintain the top-$1000$ (England) or top-$500$ (US regions) search queries (per base term). This leaves $4{,}398$ and $2{,}479$ queries respectively for England and the US regions. The second selection step is dynamic, i.e. it might have a different outcome as the training data change. In this step, we compute the linear correlation between the remaining search queries and the target variable (ILI rate) in the $5$ most recent influenza seasons in the training data (to impose recency). We maintain search queries with a correlation that exceeds a correlation threshold $\tau$, which is a learnable hyperparameter (see also Appendix D.2).
A.4 Criticism of data sets used to benchmark forecasting methods
@@
Although we present results using the ETTh1, ETTh2, and weather data sets to compare with other forecasting methods in the literature, we consider these data sets to have several shortcomings. Clearly, a weather data set that covers only a $1$-year period and is based on the sensors from just one geographical location cannot be expected to provide good enough insights into meteorological forecasting. It captures a negligible portion of meteorological data and, as a result, derived forecasting models cannot have any practical impact (compared to actual models that are used in meteorology). Hence, using this data set for training and evaluating a forecasting model may, unfortunately, result in misleading or at least inconclusive outcomes. Similar issues, but perhaps to a smaller extent, are present in the ETTh1 and ETTh2 data sets. On the one hand, these data sets seem to have a more solid practical application. However, experiments on their ETTm1 and ETTm2 variants, where the temporal resolution changes from $1$ hour to $15$ minutes (same data, shorter temporal resolution), showcase part of the problem. When we compared the forecasting performance of SOTA baselines between the two different temporal resolutions, we surprisingly found that the prediction based on the hourly sampled data set has better accuracy compared to the quarter-hourly sampled data set even at longer forecasting horizons. For example, models that conduct $192$ or even $336$ time steps (equivalent to $8$ or $14$ days) ahead forecasting with ETTh1 and ETTh2 achieve a lower MAE compared to a $720$ time steps ahead forecast with ETTm1 and ETTm2 (equivalent to $7.5$ days ahead forecast), respectively (Nie et al., 2023; Wu et al., 2023; Luo & Wang, 2024b).
@@
We also noticed that some papers run experiments on US national weekly ILI rates obtained from the CDC (Zhou et al., 2022b; Wang et al., 2023; Nie et al., 2023). The first concerning observation was that in these experiments, the input variables were based on variates (columns) from the CDC extract (spreadsheet) that only provide redundant insights compared to the ILI rate (e.g. the raw numerator and denominator of an ILI rate or the population unweighted ILI rate). In addition, some forecasting horizons that were explored in these benchmarks were practically unreasonable (even from a biological perspective, there are limits in predicting the future, especially when it comes to viruses). SOTA forecasting models for influenza generally provide good accuracy at approximately a $2$ weeks ahead forecasting horizon (Osthus & Moran, 2021; Morris et al., 2023). Anything beyond that with satisfactory performance should be considered as a very important development within the disease modelling community. We do think DeformTime might fall in that category as it delivers good results for $3$ or $4$ weeks ahead forecasting horizons. Contrary to that, in the aforementioned ILI rate benchmark, there were forecasting horizons of $24$ or $60$ weeks ahead (i.e. even more than a year!). Of course, the accuracy of the forecasting models under these forecasting horizons was reportedly detrimentally poor (influenza does not have a strong periodicity), and consequently could not be used for any meaningful comparative conclusions. If all forecasting models are performing poorly, even the least poorly performing one still is an inadequate forecaster (for the underlying task).
@@
Beyond this, we have also excluded data sets, such as traffic (Wu et al., 2021) and electricity consumption (Zhou et al., 2021), that do not have one explicitly defined target variable, but instead focus on multivariate predictions (multi-task learning). Consequently, these data sets are not entirely compatible with our forecasting task or method (DeformTime makes predictions about one target variable).
@@
Appendix B Detailed explanation in deformation using V-DAB
In this section, we provide a detailed description of how positional deformation is conducted in DeformTime, taking V-DAB as an example. Suppose we have an element $\mathbf{Z}_{\texttt{p}}(m,n)$ that was originally on position $\mathbf{p}=(m,n)$ of the patch matrix $\mathbf{Z}_{\texttt{p}}$. In this example, let’s assume that $m=5$ and $n=10$ and that the learned positional offset for that position is $(\Delta m,\Delta n)=(-0.1,1.5)$. Then, the value for $\mathbf{Z}_{\texttt{d}}(m,n)$ is the value sampled from $\mathbf{Z}_{\texttt{p}}(m+\Delta m,n+\Delta n)$ or, in this example, $\mathbf{Z}_{\texttt{p}}(5-0.1,10+1.5)$ with bilinear interpolation. Specifically, the deformed point $\mathbf{p}=(4.9,11.5)$ lies within a $2\times 2$ index matrix of actual discrete positions, i.e. $(4,11)$, $(4,12)$, $(5,11)$, and $(5,12)$. We first linearly interpolate the values across the rows of the index matrix. We note that the deformed position in this example is equal to $11.5$ and that means that there is an equal distance between the indices $11$ and $12$. We therefore perform:
$\mathbf{Z}_{\texttt{p}}(4,11.5)$ $=\mathbf{Z}_{\texttt{p}}(4,11)\times 0.5+\mathbf{Z}_{\texttt{p}}(4,12)\times(1-0.5)\,\,\text{and}$
$\mathbf{Z}_{\texttt{p}}(5,11.5)$ $=\mathbf{Z}_{\texttt{p}}(5,11)\times 0.5+\mathbf{Z}_{\texttt{p}}(5,12)\times(1-0.5)\,\,.$
We then conduct linear interpolation along the column to obtain the final value of the bilinear interpolation. Here we note that the deformed position is equal to $4.9$ and hence it does not have the same distance between indices $4$ and $5$. We therefore perform:
$\mathbf{Z}_{\texttt{p}}(4.9,11.5)=\mathbf{Z}_{\texttt{p}}(4,11.5)\times 0.1+\mathbf{Z}_{\texttt{p}}(5,11.5)\times(1-0.1)\,\,.$
This operation is conducted on every $\mathbf{p}\in\mathbf{Z}_{\texttt{p}}$ to obtain $\mathbf{Z}_{\texttt{d}}$.
$\mathbf{Z}_{\texttt{d}}$ then forms the key ($\mathbf{K}$) and value ($\mathbf{V}$) embeddings with learnable weight matrices. Given a scalar element $z_{(i,j)}\in\mathbf{Z}_{\texttt{p}}$, $\mathbf{K}$ and $\mathbf{V}$ contain elements sampled from the neighbouring latent variables ($\mathbf{z}_{i-\alpha}$, $\mathbf{z}_{i+\alpha}$) in V-DAB or neighbouring time steps $(z_{(i,j-\alpha)},z_{(i,j+\alpha)})$ in T-DAB, where $\alpha$ is the offset amplitude.
Appendix C Summary of other forecasting models